OpenStudy (awkwardpanda):

Let f(x)=x^2+5 and g(x)=x+5 / x. find (gºf)(-3).

OpenStudy (anonymous):

Is \[ g(x) =\frac{x+5}x \] ?

OpenStudy (awkwardpanda):

yes :3

OpenStudy (anonymous):

first find f(-3) \[ f(-3)=(-3)^2+5=14 \]

OpenStudy (anonymous):

find now g(14)

OpenStudy (anonymous):

Can you do that?

OpenStudy (awkwardpanda):

Not really, I'm half asleep right now. Would it be too much trouble to explain?

OpenStudy (anonymous):

\[ g(14) = \frac { 14+5}{14}=\frac {19}{14} \]

OpenStudy (anonymous):

And that is your answer.

OpenStudy (awkwardpanda):

Whoa really? thank you so much.

OpenStudy (anonymous):

YW